Determine the specific gravity of a fluid having viscosity 0.07 poise and kinematic viscosity 0.042 stokes.
1
Expert's answer
2020-11-09T19:05:40-0500

Specific gravity


"S(for\ liquids)=\dfrac{\rho_L}{\rho_{H_2O}}"

Density of water "\rho_{H_2O}=1000kg/m^3"

Given

"\mu_L=0.07\text{ poise}=7\times10^{-3}Ns/m^2,"

"\nu_L=0.042St=4.2\times10^{-6}m^2/s"

Use the relation


"\nu_L=\dfrac{\mu_L}{\rho_L}"

Then

"S=\dfrac{\rho_L}{\rho_{H_2O}}=\dfrac{\mu_L}{\nu_L\cdot\rho_{H_2O}}"

"S=\dfrac{7\times10^{-3}Ns/m^2}{4.2\times10^{-6}m^2/s\cdot1000kg/m^3}=\dfrac{5}{3}\approx1.667"